OpenStudy (anonymous):

A triangle has side lengths 10, 15, and 7. Is the triangle acute, obtuse, or right? Explain

OpenStudy (anonymous):

All I know is that its not obtuse and that i think its right.

OpenStudy (mertsj):

10^2+15^2=17^2 100+225=289 325=289 Nope. It's not a right triangle.

OpenStudy (mertsj):

If it were a right triangle, the hypotenuse would be sqrt325 or 18 Since the long side is less than 18, the triangle is acute.

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw